OpenStudy (anonymous):

Write an equation of the line containing the given point and parallel to the given line. (-2,8); 7x=5y+6

OpenStudy (shadowfiend):

If two lines are parallel, what is the relation between their slopes?

OpenStudy (anonymous):

are they equal distance?. i dunno im horrible with math, not one of my strong points

OpenStudy (shadowfiend):

It's okay :) Two parallel lines have equal slopes. So, y = 3x + 135 y = 3x + 2 Those two are parallel lines, because they the same slope (3). Let's focus on the problem below, since you'll have to use the same technique to get the equation they give you in y = mx + b form here.

OpenStudy (shadowfiend):

So now you have 7x = 5y + 6. Since we need the slope, we need to solve for y again.

OpenStudy (anonymous):

okay so would we need to subtract 6 from both sides?

OpenStudy (shadowfiend):

Yep, that's the first step.

OpenStudy (anonymous):

-6+7x=5y

OpenStudy (anonymous):

or 5y=7x-6

OpenStudy (shadowfiend):

Right. What's next?

OpenStudy (anonymous):

get the y by itself

OpenStudy (shadowfiend):

Exactly. So since the y is multiplied by 5, how do we get it by itself?

OpenStudy (anonymous):

divide

OpenStudy (shadowfiend):

Right. So we divide each part of the right by 5; what do we get?

OpenStudy (anonymous):

y=7/5x-6/5

OpenStudy (shadowfiend):

Boom! Exactly. So now we know what the slope of a parallel line would have to be. What is it?

OpenStudy (anonymous):

7/5

OpenStudy (shadowfiend):

Right. So now we'll need another equation that looks like: y = 7/5x + b Does that make sense?

OpenStudy (anonymous):

yea that does

OpenStudy (shadowfiend):

Ok, and now they've given us a point that the line goes through, so we can plug in the x and y and solve for b.

OpenStudy (anonymous):

okay so 8=7/5(-2)+b

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

Absolutely!

OpenStudy (anonymous):

So now what do I do?

OpenStudy (anonymous):

So we need to figure out what b is, right? How should we do that?

OpenStudy (anonymous):

um get b by itself?

OpenStudy (anonymous):

Exactly. So what's the first step?

OpenStudy (anonymous):

multiply 7/5(-2)

OpenStudy (anonymous):

Ok, what does that equal?

OpenStudy (anonymous):

-6/5

OpenStudy (anonymous):

Not quite. Remember you are multiplying 7/5 by -2. Do you know how to multiply fractions?

OpenStudy (anonymous):

not really is it common denominators at all?

OpenStudy (anonymous):

That is actually for adding and subtracting. Multiplication is much easier. You are multiplying 7/5 by -2. Remember that -2 is the same as -2/1, which means you have \[\frac{7}{5} \times \frac{-2}{1}\] then you just multiply straight across. What do you get?